# mO Vaskar[#1]

### v1

@@ 3 Method
Drawing on recent efforts on goal-conditioned reinforcement learning [5, 13] and representation learning [20, 31] in other areas, we show that we can learn representations that approximate a Q-function within a language reward model using contrastive learning. Since the Q-function quantifies the expected cumulative future rewards of taking a specific action from a given state [47], forcing the reward model’s output scores to be entangled with an approximated Q-function has the potential to improve the reward model’s credit assignment and preference ranking accuracy. We first present preliminaries on reward modeling and contrastive learning in Section 3.1 and present our method in Section F.
@@ 3.2 Learning Goal-Conditioned Representations in Language Reward Models
Computing Q-values. Once the reward model has been trained, we can use it to compute an additional Q-value, which gives us a measure proportional to the expected future rewards given the current state and a specific action. The Q-value at a token $t$ is the cosine similarity between the trained reward model’s hidden state, $h_{t}$, and a goal state. The the choice of goal state here is flexible. For our experiments in Section 4.1, unless specified otherwise, we set the goal state to be the mean representation of all the preferred completion in the reward model training set. Thus, we do not require any additional annotations aside from the starting preference ranking dataset. In experiments on LM steering, we construct a prototype to be used as the goal state. The prototype is constructed using generations labeled as helpful, complex, coherent, and correct by annotators in Nvidia’s HelpSteer dataset [66]. Out of the highly-scored generations, we sample a subset and use the mean represetations of the sample as the goal state for the reward model. We ablate these choices in Appendix F.
@@ 4.1.1 Experimental Setup
Our base model is OpenMath-CodeLlama-7b-Python-hf [61], a strong decoder-only LM that has been instruction tuned on OpenMathInstruct-1 to use a Python interpreter in order to solve mathematical problems, which we add a reward head to predict reward scores (Section F). This model trained with the traditional preference ranking objective serves as our baseline (Codellama 7b Reward), which we compare with the same model trained using our method (Q-Function 7b Reward).
@@ 4.1.2 Reward Modeling Evaluations
We obtain completions from the base model by greedy decoding, following prior work [61, 65], and measure the reward score AUROC. This quantifies the ability of a model’s reward score to discern between correct/incorrect sequences across different thresholds.
@@
These representations also help reward scores judge which partial solutions are promising. Part of the motivation and design of our approach is to encode the likelihood of reaching a certain goal state at intermediate steps in the generated sequence. To examine how well our method accomplishes this, we measure reward score AUROC when the model is only shown up to a certain percentage of the generated sequence (i.e., percentile). Here, we show this on the ID datasets, where we consider the 50 sample generations from the base model and plot the mean AUROC across the 50 samples. Plots for other benchmarks can be found in Appendix C. Looking at Figure 3, we find that the AUROC scores for both reward models tend to increase as they see more of the generated sequence, but Q-Function 7b Reward largely maintains higher AUROC across different percentiles and generally exhibits fewer significant drops between percentiles. These improvements across percentiles imply that the reward scores do indeed better capture which generated sequence are likely to be correct or incorrect.
@@ 4.1.3 Directly Using the Learned Representations
Figure 4: Two examples from the GSM8K dataset that was filtered via the Q-value pruning. The token level Q-values are portrayed as a heat map where the colors red and blue represents scores close to $-1$ and $1$, respectively. Both examples illustrate that the Q-values pinpoint the exact logical error in reasoning. The full version of these examples can be found in Appendix D
@@ 4.1.4 Aligning Policy Models with RLHF
We investigate whether the reward model trained with our proposed method provides a better reward signal when performing RLHF, which is an important usage of these reward models. We compare the math problem solving performance of a policy model that is aligned using a preference-ranking reward model versus our Q-Function reward model in the Proximal Policy Optimization (PPO) [58] algorithm. The prompt dataset is given by problems in the training sets of GSM8k and MATH. To introduce diversity in the prompts, which has been shown to improve PPO, we also add problems from the MathQA dataset - a large scale dataset with real world math problems [4]. We keep the settings and hyperparameters during PPO constant, and further details are in Appendix A.
Table 2 shows that, given the same settings, utilizing a reward model trained with our method improves policy model accuracy across several benchmarks. We see an average gain, weighted by benchmark size, of $1.7\%$. Intuitively, this shows that the improvements we observe in reward scores (Section 4.1.2) do translate to improvements in the policy model via RL, which further supports the value of the representations learned via our approach. Although, the gains in accuracy are not as significant compared to those observed with the reward scores. A possible explanation for this is that, due to experimental constraints such as additional data collection, the preference-ranking dataset used for reward model training are off-policy Mixtral-8x7B model  [32] generations. Hence, a clear next step for this setting would be to retrain the reward model with on-policy preference-rankings and iterate.
Model IID OOD
@@
Table 2: Accuracy of the policy model trained via PPO using a preference ranking reward model vs our Q-Function reward model. We present average accuracy across 4 independent runs. The base model results are also shown as a reference point presented in [61]. The full table with confidence intervals is in the appendix (Table 6).
@@ 4.2.2 Reward Model Representations can Steer Language Models Towards Helpful Generations
Model Helpful Correct. Coherent Complex. SFT 45.2 43.9 41.8 39.2 Prototype 54.8 56.1 58.3 60.8
@@ 5 Limitations and Future Work
Deriving more informative goal states. In Appendix F, we demonstrate the importance of having a precise and meaningful goal state. At train time, our SGS method achieves this by looking at future tokens of the preferred and dispreferred completions. During inference, however, we take the mean representation of multiple goal state completions from the training data. Although we see that the SGS method generalizes despite this discrepancy, future work could investigate other methods to derive more precise inference-time goal states.
Partial Q-values correlate with partial reward scores. An interesting observation, shown in Appendix G, is that the reward scores for partial completions are correlated with the Q-values. While our experiments show that this can benefit reward scores, especially for math reasoning, disentangling these values may be desirable in some instances. For example, one may want to incorporate the completeness of the generated sequence in the reward score, such as for creative writing, meaning incomplete sequences should strictly get low reward scores. However, currently, partial sequences may still score highly if they are likely to reach a goal state. This is expected as the final reward score is derived by a linear projection from the hidden representation, while the Q-values are derived as the cosine similarity between that same representation and goal state. In Appendix G, we explore the effects of learning an MLP projection for the rewards to help decouple the two signals. Further disentangling these signals in the representations could be an interesting direction for future work.
@@ A.3 RLHF Training and Evaluation Details
Table 6: Average accuracy and $95\%$ confidene interval after the base policy model was trained via the PPO algorithm using a preference ranking reward model and our Q-Function reward model, across 4 different independent runs.
@@ B.3 Compute and Statistical Significance
Appendix C Partial Completion Rewards
Appendix D Q-Value Visualization
D.1 Good Examples
D.2 Noisy Examples
Appendix E Exploring Sources of Goal States
@@
Appendix F Source State and Goal State Sampling Ablation
In this section, we discuss ablations to source and goal state sampling. In computing the contrastive loss during training, we need to sample a positive source and goal state pair as well as a negative source and goal state pair. To investigate how we should sample these tokens, we explore 3 different approaches.
@@ Appendix F Source State and Goal State Sampling Ablation
Appendix G Reward Projection Ablation
In this section, we show high correlation between our learned Q-values and reward scores for partial completions, as illustrated in Table 10. The partial completions are from the base model on the GSM8K test split [16]. This is expected since, by design, the reward model outputs are dependent on the learned features that approximate the Q-function. This presents a potential risk in practice that the policy model may exploit during PPO. We explore attempting to decouple the reward score from the Q-value for partial completions by projecting the hidden features which approximate the Q-value through a multilayer perceptron. Table 10 and Table 11 compares the result of projecting the hidden features via a MLP layer versus a linear layer. The finding is we can reduce correlation between rewards and Q-values for partial completions while maintaining high reward model performance. Iterative retraining of the reward model on on-policy completions would further mitigate any potential reward model hacking given that our complete loss function also considers the preference ranking paired loss in 1.
@@ Appendix G Reward Projection Ablation
Appendix H Examples of Prototype Guided Decoding

### v2

# Learning Goal-Conditioned Representations for Language Reward Models

@@ 3 Method
Drawing on recent efforts on goal-conditioned reinforcement learning [5, 13] and representation learning [20, 31] in other areas, we show that we can learn representations that approximate a Q-function within a language reward model using contrastive learning. Since the Q-function quantifies the expected cumulative future rewards of taking a specific action from a given state [47], forcing the reward model’s output scores to be entangled with an approximated Q-function has the potential to improve the reward model’s credit assignment and preference ranking accuracy. We first present preliminaries on reward modeling and contrastive learning in Section 3.1 and present our method in Section D.2.
@@ 3.2 Learning Goal-Conditioned Representations in Language Reward Models
Computing Q-values. Once the reward model has been trained, we can use it to compute an additional Q-value, which gives us a measure proportional to the expected future rewards given the current state and a specific action. The Q-value at a token $t$ is the cosine similarity between the trained reward model’s hidden state, $h_{t}$, and a goal state, $h_{g}$. The the choice of goal state here is flexible. For our experiments in Section 4.1, unless specified otherwise, we set the goal state to be the mean representation of all the preferred completion in the reward model training set. More explicity, the goal state is computed as:
$h_{g}=\frac{1}{N}\sum_{y_{w}\in\mathcal{D}}h({y_{w}})$ (7)
where $h(.)$ represents the reward model’s last hidden layer of the final token of the completion. In Appendix C, we ablate this choice and find this to be the best method for computing the goal state during inference. Thus, we do not require any additional annotations aside from the starting preference ranking dataset. In experiments on LM steering, we construct a prototype to be used as the goal state. The prototype is constructed using generations labeled as helpful, complex, coherent, and correct by annotators in Nvidia’s HelpSteer dataset [67]. Out of the highly-scored generations, we sample a subset and use the mean represetations of the sample as the goal state for the reward model. We ablate these choices in Appendix D.2.
@@ 4.1.1 Experimental Setup
Our base model is OpenMath-CodeLlama-7b-Python-hf [62], a strong decoder-only LM that has been instruction tuned on OpenMathInstruct-1 to use a Python interpreter in order to solve mathematical problems, which we add a reward head to predict reward scores (Section D.2). This model trained with the traditional preference ranking objective serves as our baseline (Codellama 7b Reward), which we compare with the same model trained using our method (Q-Function 7b Reward).
@@ 4.1.2 Reward Modeling Evaluations
We obtain completions from the base model by greedy decoding, following prior work [62, 66], and measure the reward score AUROC. This quantifies the ability of a model’s reward score to discern between correct/incorrect sequences across different thresholds. In Appendix E, we additionally assess Q-Function 7b Reward’s ability to rank completions.
@@
These representations also help reward scores judge which partial solutions are promising. Part of the motivation and design of our approach is to encode the likelihood of reaching a certain goal state at intermediate steps in the generated sequence. To examine how well our method accomplishes this, we measure reward score AUROC when the model is only shown up to a certain percentage of the generated sequence (i.e., percentile). Here, we show this on the ID datasets, where we consider the 50 sample generations from the base model and plot the mean AUROC across the 50 samples. Plots for other benchmarks can be found in Appendix F. Looking at Figure 3, we find that the AUROC scores for both reward models tend to increase as they see more of the generated sequence, but Q-Function 7b Reward largely maintains higher AUROC across different percentiles and generally exhibits fewer significant drops between percentiles. These improvements across percentiles imply that the reward scores do indeed better capture which generated sequence are likely to be correct or incorrect.
@@ 4.1.3 Directly Using the Learned Representations
Figure 4: Two examples from the GSM8K dataset that was filtered via the Q-value pruning. The token level Q-values are portrayed as a heat map where the colors red and blue represents scores close to $-1$ and $1$, respectively. Both examples illustrate that the Q-values pinpoint the exact logical error in reasoning. The full version of these examples can be found in Appendix G
@@ 4.1.4 Aligning Policy Models with RLHF
We investigate whether the reward model trained with our proposed method provides a better reward signal when performing RLHF, which is an important usage of these reward models. We compare the math problem solving performance of a policy model that is aligned using a preference-ranking reward model versus our Q-Function reward model in the Proximal Policy Optimization (PPO) [59] algorithm. We additionally compare the performance against DPO as it serves as a strong baseline  [55]. The prompt dataset is given by problems in the training sets of GSM8k and MATH. To introduce diversity in the prompts, which has been shown to improve PPO, we also add problems from the MathQA dataset - a large scale dataset with real world math problems [4]. We keep the settings and hyperparameters during PPO constant, and further details are in Appendix A.
Table 2 shows that, given the same settings, utilizing a reward model trained with our method improves policy model accuracy across several benchmarks. We see an average gain, weighted by benchmark size, of $1.7\%$. Intuitively, this shows that the improvements we observe in reward scores (Section 4.1.2) do translate to improvements in the policy model via RL, which further supports the value of the representations learned via our approach. Although, the gains in accuracy are not as significant compared to those observed with the reward scores. A possible explanation for this is that, due to experimental constraints such as additional data collection, the preference-ranking dataset used for reward model training are off-policy Mixtral-8x7B model  [32] generations. Hence, a clear next step for this setting would be to retrain the reward model with on-policy preference-rankings and iterate. Interestingly, we observe that DPO performs on par with our method for the ID tasks (GSM8k and MATH). However, the policy trained with the contrastive RM performs much better on OOD tasks (algebra222, GSM-Hard, Asdiv, mawps, and svamp) when compared to DPO. These findings could be explained by related works that analyze DPO and find it can perform poorly in OOD settings  [72]. All in all, these results indicate that PPO training with the contrastive RM leads to better generalization than training with DPO, particularly in OOD settings.
Model ID OOD
@@
Table 2: Accuracy of the policy model trained via PPO using a preference ranking reward model vs our Q-Function reward model. We present average accuracy across 4 independent runs. The base model results are also shown as a reference point presented in [62] and DPO results are shown as a strong baseline. The full table with confidence intervals is in the appendix (Table 6).
@@ 4.2.2 Reward Model Representations can Steer Language Models Towards Helpful Generations
Model Helpful Correct. Coherent Complex.
SFT 45.2 43.9 41.8 39.2
Prototype 54.8 56.1 58.3 60.8
@@ 5 Limitations and Future Work
Deriving more informative goal states. In Appendix D.2, we demonstrate the importance of having a precise and meaningful goal state. At train time, our SGS method achieves this by looking at future tokens of the preferred and dispreferred completions. During inference, however, we take the mean representation of multiple goal state completions from the training data. Although we see that the SGS method generalizes despite this discrepancy, future work could investigate other methods to derive more precise inference-time goal states.
Partial Q-values correlate with partial reward scores. An interesting observation, shown in Appendix H, is that the reward scores for partial completions are correlated with the Q-values. While our experiments show that this can benefit reward scores, especially for math reasoning, disentangling these values may be desirable in some instances. For example, one may want to incorporate the completeness of the generated sequence in the reward score, such as for creative writing, meaning incomplete sequences should strictly get low reward scores. However, currently, partial sequences may still score highly if they are likely to reach a goal state. This is expected as the final reward score is derived by a linear projection from the hidden representation, while the Q-values are derived as the cosine similarity between that same representation and goal state. In Appendix H, we explore the effects of learning an MLP projection for the rewards to help decouple the two signals. Further disentangling these signals in the representations could be an interesting direction for future work.
@@ A.3 RLHF Training and Evaluation Details
Table 6: Average accuracy and $95\%$ confidene interval after the base policy model was trained via DPO, PPO using a preference ranking reward model, and PPO using our Q-Function reward model, across 4 different independent runs.
@@ B.3 Compute and Statistical Significance
Appendix C Inference-Time Goal State Ablation
In this appendix section, we explore different choices of inference-time goal states and compare their performance. To ablate this, we conduct a new experiment under the mathematical reasoning with code setting (Section 4.1) where we evaluate the effect of (a) choosing poor examples to construct the goal state and (b) the number of examples used to construct the goal state. In particular, we evaluate 3 settings for picking the sample completions for constructing the goal state:
  1. 1.
    We vary the number of preferred completions used to construct the goal state to evaluate the effect of adding more examples of good reasoning.
  2. 2.
    We vary the number of dispreferred completions used to construct the goal state, to evaluate whether bad examples lead to poor performance.
  3. 3.
    We incrementally add more dispreferred completions to a fixed sample of preferred completion in order to measure the robustness of our goal state computation to negative examples. We refer to the addition of dispreferred completions on top of all preferred completions as adding “corrupted” examples.
The results are in Figure 5. Overall, these results show that negative or unhelpful completions degrade performance, indicating the importance of choosing relevant examples for computing the goal state. Additionally, they demonstrate that having more examples of the concept leads to better performance as more generations are filtered with comparable proportion of those remaining generations being correct.
(a) MATH
(b) GSM8K
Figure 5: We plot the average proportion of non-filtered responses that are correct and the average number of completions (out of a total 50) that are filtered, as we vary the sample size (1, 10, 100, 1K, 10K, and 30K) of dis-preferred, preferred, and corrupted examples when constructing the goal state. All of the examples are retrieved from the preference ranking training dataset. We plot the performance metrics for the MATH and GSM8K test sets. Each of the test sets has 50 completions per problem that is generated by the base model and we filter any completion for which the Q-value for any token in the completion sequence is below 0. At each sample size, we run 3 independent random sample from the full set of preferred and dis-preferred completions to construct 3 independent goal states, with 95% CI.
We additionally ablate our choice of using the last token in the completion sequence to construct the goal state. In this experiment, we repeat our filtering experiments from Section 4.1.3. Except, we randomly sample a token from in the completion sequence and use it as the goal state:
Sampling Method GSM8k MATH
Accuracy (%) Prop. Correct (%) Accuracy (%) Prop. Correct (%)
Last Token 86.0 84.0 59.6 52.0
Random Token 85.9 81.8 57.3 45.3
Table 8: Majority vote accuracy and average proportion correct when filtering examples using last token and random token sampling for constructing the goal state.
From the result of this ablation, we see that last token sampling has superior performance, likely because the last token on average carries the most information about what good & complete generations look like.
Appendix D Train-Time Goal State Ablation
D.1 Exploring Sources of Goal States
@@
D.2 Source State and Goal State Sampling Ablation
In this section, we ablate the sampling strategy for the source and goal states. In computing the contrastive loss during training, we need to sample a positive source and goal state pair as well as a negative source and goal state pair. To investigate how we should sample these tokens, we explore 3 different approaches.
@@ D.2 Source State and Goal State Sampling Ablation
Appendix E Reward Model Ranking Performance
In this appendix section, we compare the Q-Function 7b Reward to Codellama 7b Reward by judging their respective abilities to rank completions. Since we have a total of 50 generations per problem, we use the baseline reward mode and our reward model to select the Top 1, 5, 10, and 25 samples as ranked by the reward scores. With the selected sample, we perform majority vote and also note the average proportion of the sample K that are correct solutions. The results are provided below. We show the Majority Vote Accuracy and Average Proportion Correct when selecting Top-K completions in order of reward model scores. The two reward models used here are the baseline Codellama RM and our Q-Function RM.
GSM8k MATH
Model Top-K Accuracy (%) Prop. Correct (%) Accuracy (%) Prop. Correct (%)
1 84.6 84.6 51.7 51.7
Q-Function RM 5 86.2 84.5 59.7 51.1
10 86.2 83.8 59.5 50.3
25 85.5 81.8 57.8 47.0
1 80.8 80.8 43.8 43.8
Codellama RM 5 84.3 81.9 54.0 45.8
10 85.2 82.1 56.0 46.9
25 85.8 81.1 56.5 46.1
Table 11: Majority Vote Accuracy and Average Proportion Correct when selecting Top-K completions in order of reward model scores. The two reward model use here is the baseline Codellama RM and our Q-Function RM.
The results show that our Q-Function RM clearly outperforms the baseline.
Appendix F Partial Completion Rewards
Appendix G Q-Value Visualization
G.1 Good Examples
G.2 Noisy Examples
Appendix H Exploring Correlation Between Q-values and Reward Scores
In this section, we explore the correlation between our learned Q-values and reward scores for partial completions, as illustrated in Table 12, and empirically investigate the possibility of reward hacking during PPO training. The partial completions are from the base model on the GSM8K test split [16]. The correlation is expected since, by design, the reward model outputs are dependent on the learned features that approximate the Q-function. This presents a potential risk in practice that the policy model may exploit during PPO. We explore attempting to decouple the reward score from the Q-value for partial completions by projecting the hidden features which approximate the Q-value through a multilayer perceptron. Table 12 and Table 13 compares the result of projecting the hidden features via a MLP layer versus a linear layer. The finding is we can reduce correlation between rewards and Q-values for partial completions while maintaining high reward model performance. Iterative retraining of the reward model on on-policy completions would further mitigate any potential reward model hacking given that our complete loss function also considers the preference ranking paired loss in 1.
@@ Appendix H Exploring Correlation Between Q-values and Reward Scores
To empirically evaluate the risk of the policy gaming the reward function, we conduct a new experiment where we run PPO training for an extended time (5 episodes) in order to evaluate whether reward hacking occurs in practice. The reward scores and response lengths are shown in Figure 7 and the performance of the model at the end of each episode is given in Table 14.
Figure 7: Train time average rewards and average response length for PPO using Codellama RM (purple) and PPO using our Q-Function RM (brown)
Model Episode GSM8k MATH algebra222 GSM-Hard Asdiv mawps svamp
1 79.7 43.6 66.3 61.2 77.9 92.0 78.5
2 79.6 45.0 72.5 61.5 78.0 91.6 79.9
Codellama 3 80.5 45.3 69.8 62.0 78.3 92.3 80.9
4 80.5 45.5 70.7 62.5 78.3 92.0 79.5
5 80.5 45.2 70.7 62.5 78.6 92.0 80.5
1 80.2 45.9 67.6 62.0 79.2 93.9 81.6
2 80.6 46.5 74.3 61.9 79.5 93.1 80.3
Q-Function 3 81.7 46.6 73.0 63.5 79.2 93.2 81.6
4 81.0 46.5 73.4 63.2 79.2 93.4 81.0
5 81.1 46.5 74.3 63.0 79.8 93.6 81.0
Table 14: PPO-trained policy model accuracy across several math benchmarks comparing using the baseline Codellama RM and our Q-Function RM for a single long 5 episode training run.
These results indicate training is fairly stable and the policy model does not game the reward function in practice. One explanation for why gaming does not occur is that typically during RLHF a relatively low learning rate is used, making large shifts in the models generations less likely to occur.
Appendix I Examples of Prototype Guided Decoding